OpenStudy (rianiscool):

need help asap will give a medal Enter the answer. 2. Evaluate. (34 + 14) ÷ [42 – (20 + 16)]

OpenStudy (anonymous):

simplified it equals 8

OpenStudy (lucidiousxiv):

Basic Pemdas. Complete the work in the brackets first, then do both sets of parenthesis, and then basic math. It all works out to equal 8. 48 ÷ (42 - 36) = 48 ÷ 6 Answer = 8

OpenStudy (anonymous):

ha i got it first :P

OpenStudy (lucidiousxiv):

That is supposed to be this. \[48\div(42-36)\] then \[48\div6=8\]

OpenStudy (anonymous):

but i give you medal anyway :D

OpenStudy (rianiscool):

what

OpenStudy (anonymous):

the answer is 8

OpenStudy (rianiscool):

thank you all

OpenStudy (anonymous):

np

OpenStudy (lucidiousxiv):

Pemdas = Parenthesis Exponents Multiplication Division Addition Subtraction Which is what type of method you would have to follow in this current problem.

OpenStudy (rianiscool):

:)